# A Benchmark for Data Management in Microservices

### v2

@@ 7.4. A Composite Solution
We implement two replication semantics: (i) Eventual, enabled by Orleans Streams (Orleans, 2024b), an Orleans add-on that allows actors to subscribe and publish to streams dynamically; and (ii) Causality across multiple objects, enabled by a primary-secondary Redis deployment. To decouple writers from readers, we use the following scheme: Product actors write to the primary node, and updates are asynchronously streamed to the secondary node, so Cart actors can read from it. As Redis linearizes writes, updates are applied to the Cart in order.
8. Example Experiments
In this section, we show example results of our Online Marketplace implementations in Orleans and Statefun to investigate the characteristics of our benchmark under different concurrency and scaling levels, and workload skewness . The example experiments on scalability are run on a multi-core machine, which is also sufficient for many applications. Extensive experiments on a large cluster need substantial efforts that need to be addressed in future work.
8.1. Experimental Settings
Deployment. We set up our benchmarking environment on UCloud (UCloud, [n.d.]b) based on u1-standard instances. A u1-standard contains an Intel Xeon Gold 6130 CPU@2.10 GHz, 32 vCPUs, and 384 GB of memory (UCloud, [n.d.]a). UCloud instances run inside a Kubernetes cluster connected via 100Gbps Infiniband virtual network.
To avoid resource competition, we allocate independent instances to different benchmarking components, namely, the benchmark driver, the target platform, the PostgreSQL database server, and the Redis instances. To remove cache effects, we restart Orleans and Statefun after each run. In experiments involving Redis and PostgreSQL, we also clean up the state after each run. Besides, after data population, we ensure CPU usage returns to idle before initializing transaction submission. We use PostgreSQL 14.5 running in the operating system (OS) Debian 12.1, whereas the driver and platform instances run in the OS Ubuntu 22.10. All the instances are located in the same region and availability zone.
@@ 8.1. Experimental Settings
Workload and Parameters. Table 2 lists the benchmark driver input parameters used in the experiments. We set the probability of payments accepted and checkout to 100% to maximize the amount of transactions. The defined transaction ratio targets mimicking the popular scenario where customer orders contain products from multiple sellers, thus requiring multiple deliveries per order. We use both uniform and skewed distribution for picking sellers per each transaction. The number of sellers and products are picked so to not introduce substantial conflicts in uniform distribution.
Platform Configuration. Statefun introduces a larger configuration space compared to Orleans. To mitigate the possible effects of suboptimal parameters, we systematically experimented many different parameters until we reached the ones that yield the optimal performance under our experimental setup. These parameters can be found in our extended version (Laigner et al., 2023).
Table 2. Experiment Parameters
Configuration Parameter Value
Data Population Customers 100K
Products 100K
Sellers 10K
Transaction Ratio Update Delivery 60%
Customer Checkout 30%
Seller Dashboard 5%
Update Price 3%
Delete Product 2%
Customer Checkout Max Number of Cart Items 10
Checkout Probability 100%
Voucher Probability 5%
Distribution Seller UNIFORM/ZIPFIAN
Product UNIFORM
8.2. Experiment Results
Driver Scalability. To examine if the safety mechanisms of the driver (Section 6.2) are scalable, we conduct a scalability experiment. We use the parameters shown in Table 2 and fix the transaction latency to 100 ms, varying the number of concurrent workers submitting transactions. We observe in Figure 2(a) that the number of completed transactions scales with the concurrency level, showing stable throughput across epochs. The overhead entailed by adding more worker threads does not create a bottleneck, independent of the distribution picked for sellers and products.
@@ 8.2. Experiment Results
As shown in Figure 2(a), Orleans achieves maximum throughput with the concurrency level set to 48. With an even higher concurrency level, more messages would be accumulated on actors’ input queue, particularly actors in the checkout critical path (e.g., with Order and Payment doubling their processing latency).
The overhead introduced by logging is negligible at the start and remains constant as the concurrency level increases, achieving a maximum of 12% overhead compared to non-logging Orleans. The lower throughput is explained by the wait introduced by write requests submitted to PostgreSQL.
As for Statefun (Figure 2(b)), in order to achieve maximum throughput, the concurrency level required is significantly higher compared to Orleans. We conjecture several reasons: (i) Although we aimed for the most performant deployment style, the containerized deployment mode introduces overhead on execution functions due to the virtualization layer;  [^1] We could not confirm this overhead because we do not find instructions in the documentation to run Statefun in bare metal. (ii) The HTTP ingress acts as a single operator, processing requests serially, introducing a bottleneck;  [^2] We tried using RichParallelSourceFunction to enable parallel ingress, but we found that driver requests were arriving out of order in functions, impeding the execution of some transactions. (iii) Although we made sure to adjust the polling rate to a configuration that maximizes Statefun performance, the polling for transaction results coming from the driver to compute metrics in a timely manner invariably introduces a processing overhead.
@@
48163222224444666688881010101012121212(a) Number of CPUs on OrleansEnd-to-end Latency (milliseconds)Update DeliveryCheckoutSeller DashboardPrice UpdateProduct Delete
48163250505050100100100100150150150150200200200200250250250250(b) Number of CPUs on Statefun
@@
Preliminary experiments showed that fixing seller distribution and varying product skewness showed similar throughput across different skew levels, which led us to investigate the effects of seller skewness. Thus, we picked the sellers using a Zipfian distribution while picking products using a uniform distribution. Again, we use 32 CPUs and the concurrency level that maximizes throughput. Figure 3(c) exhibits the Zipfian value of six skew levels used.  [^3] Statefun crashes on skew levels 1 and 1.2. We show the values captured up to the crash.
It is observed that the throughput of Orleans decreases with increasing skewness, following a stable trend. This phenomenon is expected since there is a high contention on certain sellers and their products. Statefun, on the other hand, is less sensitive to workload skewness. In this case, we conjecture that the effects of batching messages play a role, allowing the functions not to incur overhead as skewness increases.
@@
Figure 4(a) shows Orleans scales linearly as we increase the number of CPUs. In a similar way, Orleans with logging scales nearly linearly. The effects of logging remain low at the start, slightly increasing from 16 CPUs on and remaining stable afterward, matching the expected overhead found in concurrency experiments (Figure 2(a)). Statefun can also take advantage of increased computational resources but to a lower degree compared to Orleans. The overhead of logging is nonexistent at the start, and from 16 CPUs on, it impacts Statefun scalability.
@@
As for Statefun, for customer checkout and update delivery, it is possible to observe latency significantly decreases as more CPUs are available, showing improvements up to 61% from 4 to 16-32. On the other hand, the overall throughput results are reflected in the longer end-to-end latency. One of the aspects that drives Statefun’s overall latency higher compared to Orleans is inherent to the programming model. Whenever an operation involves interacting with multiple functions, a function must send a message to each function and wait for the eventual arrival of the responses. The function must keep track of the responses received through custom-made code, which necessarily involves storing responses in the function’s state. On the other hand, in Orleans, multiple actors’ calls are encapsulated through promises and do not involve state operations.
@@
Turning our attention to the composite solution, Figure  4(b) shows the overhead of consistent query (+View) and different replication semantics (Eventual and Causal). The results show that the overhead introduced by causal replication is not substantial compared to eventual, which can motivate users to overlook weak semantics in their deployments. On the other hand, the overhead of replication tends to increase as more transactions compete for resources. In the eventual case, we conjecture that, as Orleans Streams are also implemented as actors (Orleans, 2024a), they compete with transactional actors for scheduling turns. Similarly, in the causal case, the cost of the increasing number of network I/O operations tends to overcome the benefits of memory-resident data provided by Redis. We omit the results with logging because all run with an average overhead of 5 to 10%.
@@
It is worth noting that we managed to implement the missing criteria adding less than 15% overhead on average, demonstrating that meeting all the benchmark criteria is realistic. However, the composite solution performance is bounded by Orleans Transactions. Recent research shows it is possible to speed up Orleans Transactions by 4X through deterministic scheduling (Liu et al., 2022). Another direction to optimize the performance of Online Marketplace transactions is by providing weaker isolation levels that still fulfill the constraints. For instance, customer statistics can be updated in any order as long as the events are processed exactly once.
9. Related Work
TPC-W (Council, 2005) is a transactional benchmark that models the core aspects of user experience on an e-commerce website, such as browsing pages, checking out books, and searching for keywords (e.g., by title). TPC-C (Council, 2010) was designed to represent the transaction processing requirements of a wholesale supplier. YCSB (Cooper et al., 2010) models transactional workloads in the cloud that are not necessarily executed under ACID semantics. All of them assume a traditional monolithic architecture. Our benchmark models a modern microservice-oriented application and the complex interplay of their components through events, differing substantially from these benchmarks in terms of architectural style and data management requirements.
10. Conclusion
Online Marketplace was designed to incorporate data management requirements and challenges faced by microservice practitioners. Through implementing Online Marketplace in state-of-the-art platforms for developing distributed stateful applications, we encountered several shortcomings that prevented us from fulfilling the prescribed data management criteria, indicating that Online Marketplace can effectively pinpoint the missing core data management features pursued by practitioners. As a result, Online Marketplace will support designing futuristic data management systems for microservices.

### v4

@@ 7.1.4. A Composite Solution
We implement two replication semantics: (i) Eventual, enabled by Orleans Streams (Orleans, 2024c), an Orleans add-on that allows actors to subscribe and publish to streams dynamically; and (ii) Causality across multiple objects, enabled by a primary-secondary Redis deployment. To decouple writers from readers, we use the following scheme: Product actors write to the primary node, and updates are asynchronously streamed to the secondary node, so Cart actors can read from it. As Redis linearizes writes, updates are applied to the Cart in order.
L#4. Achieving consistent replication and querying requires weaving together heterogeneous systems through non-native APIs.
7.2. Experimental Study
7.2.1. Experimental Settings
Deployment. We set up our benchmarking environment on UCloud (UCloud, [n.d.]b) based on u1-standard instances. A u1-standard contains an Intel Xeon Gold 6130 CPU@2.10 GHz, 32 CPUs, and 384 GB of memory (UCloud, [n.d.]a). UCloud instances run inside a Kubernetes cluster connected via 100Gbps Infiniband virtual network. To avoid resource competition, we allocate independent instances to different benchmarking components, namely, the benchmark driver, the target platform, the PostgreSQL database server, and the Redis instances. To remove cache effects, we restart Orleans and Statefun after each run. In experiments involving Redis and PostgreSQL, we also clean up the state after each run. Besides, after data population, we ensure CPU usage returns to idle before initializing transaction submission. We use PostgreSQL 14.5 running in the operating system (OS) Debian 12.1, whereas the driver and platform instances run in the OS Ubuntu 22.10. All the instances are located in the same region and availability zone. The criteria in §$\S$4.6 are only applied to multi-node deployments. Since we focused our investigation on the performance and overhead of the platforms in multi-core settings, these are not applicable.
@@ 7.2.1. Experimental Settings
Workload and Parameters. We use the workload modeling scenario A (§$\S$ 5). We set the probability of payments accepted and checkout to 100% to maximize the amount of transactions. Checkouts contain up to 10 items and quantities are selected randomly for each item (1-5). Each cart item has 5% probability of being assigned a discount. The discount value is randomly picked, not surpassing 10% of the item’s price. We use both uniform and skewed distribution for picking sellers per each transaction. The number of sellers and products are picked reflecting size factor 10, so to not introduce substantial conflicts in uniform distribution.
7.2.2. Experiment Results
Driver Scalability. To examine if the coordination mechanisms of the driver (Section 6.2) are scalable, we conduct a scalability experiment. We use the parameters of scenario A (§$\S$ 5) and fix the transaction latency to 100 ms, varying the number of concurrent workers submitting transactions. We observe in Figure 2(a) that the number of completed transactions scales with the concurrency level, showing stable throughput across epochs. The overhead entailed by adding more worker threads does not create a bottleneck, independent of the distribution picked for sellers and products, reflecting the efficiency of the driver’s coordination mechanisms.
@@ 7.2.2. Experiment Results
As shown in Figure 2(b), Orleans achieves maximum throughput with the concurrency level 48. With an even higher concurrency level, more messages would be accumulated on actors’ input queue, particularly actors in the checkout critical path (e.g., with Order and Payment doubling their processing latency). The overhead introduced by logging is negligible at the start and remains constant as the concurrency level increases, achieving a maximum of 12% overhead compared to non-logging. The lower throughput derives from the wait introduced by updates submitted to PostgreSQL.
As for Statefun (Figure 2(c)), in order to achieve maximum throughput, the concurrency level required is significantly higher compared to Orleans. We conjecture several reasons: (i) Although we aimed for the most performant deployment style, the containerized deployment mode introduces overhead on execution functions due to the virtualization layer;  [^1] We could not confirm this overhead because we do not find instructions in the documentation to run Statefun in bare metal. (ii) The HTTP ingress acts as a single operator, processing requests serially, introducing a bottleneck;  [^2] We tried using RichParallelSourceFunction to enable parallel ingress, but we found that driver requests were arriving out of order in functions, impeding the execution of some transactions. (iii) Although we made sure to adjust the polling rate to a configuration that maximizes Statefun performance, the polling for transaction results coming from the driver to compute metrics in a timely manner invariably introduces a processing overhead.
@@
L#5. The inherent deployment and processing model of the dataflow architecture of Statefun are key factors that harm performance in transactional and event-based workloads.
48163222224444666688881010101012121212(a) Number of CPUs on OrleansEnd-to-end Latency (milliseconds)
48163250505050100100100100150150150150200200200200250250250250(b) Number of CPUs on StatefunUpdate DeliveryCustomer CheckoutSeller DashboardPrice UpdateProduct Delete
@@
Preliminary experiments showed that fixing seller distribution and varying product skewness showed similar throughput across different skew levels, which led us to investigate the effects of seller skewness. Thus, we picked the sellers using a Zipfian distribution while picking products using a uniform distribution. Again, we use 32 CPUs and the concurrency level that maximizes throughput. Figure 3(c) exhibits the Zipfian value of six skew levels used.  [^3] Statefun crashes on skew levels 1 and 1.2. The metrics are captured up to the crash.
It is observed that the throughput of Orleans decreases with increasing skewness, following a stable trend. This phenomenon is expected since there is a high contention on certain sellers and their products. Statefun, on the other hand, is less sensitive to workload skewness. In this case, we conjecture that the effects of batching messages play a role, allowing the scheduling of functions not to incur overhead as skewness increases.
L#6. A skewed execution introduces a lower performance penalty in Statefun given the inherent batch-oriented execution model.
@@
Figure 4(a) shows Orleans scales linearly as we increase the number of CPUs. In a similar way, Orleans with logging scales nearly linearly. The effects of logging remain low at the start, slightly increasing from 16 CPUs on and remaining stable afterward, matching the expected overhead found in concurrency experiments (Figure 2(b)). Statefun can also take advantage of increased computational resources but to a lower degree compared to Orleans. The overhead of logging is nonexistent at the start, and from 16 CPUs on, it impacts Statefun scalability.
@@
L#7. Ensuring all-or-nothing atomicity and concurrency control appear as dominant overheads on both platforms.
@@
As for Statefun, for customer checkout and update delivery, it is possible to observe latency significantly decreases as more CPUs are available, showing improvements up to 61% from 4 to 16-32. On the other hand, the overall throughput results are reflected in the longer end-to-end latency. One of the aspects that drives Statefun’s overall latency higher compared to Orleans is inherent to the programming model. Whenever an operation involves interacting with multiple functions, a requester function must send a message to each recipient function and wait for the eventual arrival of the responses. The function must keep track of the responses received through custom-made code, which necessarily involves storing responses in the function’s state. On the other hand, in Orleans, multiple actors’ calls are encapsulated through promises and do not involve state operations. This scenario occurs in both checkout and update delivery transactions to meet properties 1, 3, 6, and 7.
L#8.The absence of native primitives for coordinating asynchronous operations across multiple functions significantly impacts Statefun’s overall performance.
@@
Turning our attention to the composite solution, Figure  4(b) shows the overhead of consistent query (+View) and different replication semantics (Eventual and Causal). The results show that the overhead introduced by causal replication is not substantial compared to eventual, which can motivate users to overlook weak semantics in their deployments. On the other hand, the overhead of replication tends to increase as more transactions compete for resources. In the eventual case, we conjecture that, as Orleans Streams are also implemented as actors (Orleans, 2024b), they compete with transactional actors for scheduling turns. Similarly, in the causal case, the cost of the increasing number of network I/O operations tends to overcome the benefits of memory-resident data provided by Redis. We omit the results with logging because all run with an average overhead of 5 to 10%. It is worth noting that we managed to implement the missing criteria adding less than 15% overhead on average, demonstrating that meeting all the benchmark criteria is realistic. However, the composite solution performance is bounded by Orleans Transactions.
L#9. Offloading replication and continuous queries to external systems inevitably leads to increased network I/O overhead.
7.2.3. Design Decisions
In this section, based on the observations and lessons learned identified through our case studies, we devise a set of design decisions to advance data platforms and foment new ones.
In the feature space, we posit the following to address L#1, L#4, and L#9:
D#1. Enhancing microservice platforms with native, system-level support for data replication and querying microservice states is a promising feature. This must be paired with performant system-level implementations. The inherent design of existing data platforms may limit the embracement of these missing features, leading to the next two design decisions.
D#2. Integrating external systems and their features into microservice platforms is another potential candidate to relieve the burden on developers. The case study showed it is realistic to rely on features provided by external systems. However, fulfilling missing requirements must not jeopardize performance.
D#3. Integrating the features from external systems must be accompanied by optimizations to hide latencies associated with cross-system coordination to prevent substantial performance penalty.
In terms of performance, lessons L#5 and L#8 demonstrate the values of the following design decisions:
D#4. The event processing model and task scheduling in microservice platforms should reduce the impact of the long latency of certain multi-microservice workflows (e.g., customer checkout) on shorter transactions (e.g., price updates). These facilities should ideally be transparent to developers, as in Orleans. In addition, designing primitives to circumvent concurrency limitations inherent to the programming model is another opportunity, as follows.
D#5. Multiplexing multiple non-conflicting transactions within a scheduling unit (e.g., enabling grain reentrancy in Orleans) is a promising technique to speed up microservice workloads, enhancing throughput by running non-conflicting transactions concurrently.
From a system architecture perspective, lessons L#2 and L#6 derive the following:
D#6. Holistically managing state updates and message queue operations found in Statefun appears to be a promising approach for ensuring exactly-once semantics. That relieves developers from writing idempotent code to prevent duplicate executions, which may not always be possible. These should be paired with efficient solutions to schedule application functions.
D#7. Batching multiple conflicting transactions is an effective performance technique for highly-contended microservice workloads. Statefun inherits the operator execution model of Flink, which processes event streams in batches. Orleans, on the other hand, tracks message timeouts individually, which limits batching opportunities.
Lesson L#7 derives better algorithmic design and lesson L#3 addresses the usefulness of proper abstractions, respectively:
@@ 7.2.3. Design Decisions
D#8. Devising new algorithms and techniques for achieving all-or-nothing atomicity and concurrency control is critical for microservice performance. Recent research shows it is possible to speed up Orleans Transactions by 4X through deterministic scheduling (Liu et al., 2022). Offering weaker isolation levels that can still achieve the required properties is another promising direction. For instance, customer statistics can be updated in any order as long as the events are processed exactly once.
D#9. Virtualized representation of application functions and associated states, such as through a virtual actor in Orleans, appears as an effective abstraction to manage deployment, discovery, distribution, scheduling, and failures holistically in microservices.
8. Concluding Remarks
Our experiments with five implementations of Online Marketplace in three competing systems have shown varied results under different workloads, showing the ability of Online Marketplace to pinpoint missing core data management features and the impact of the programming abstractions and architectural design of these systems on performance. The experiments also highlight the overhead of all-or-nothing atomicity, concurrency control, constraint enforcement, replication semantics, consistent queries, audit logging, and asynchronous processing of events. It is worth noting these cannot be obtained by existing microservice benchmarks (§$\S$ 2). Furthermore, from the lessons learned obtained through Online Marketplace, we derive design decisions to advance the development of data systems, highlighting that Online Marketplace will support designing futuristic data management systems for microservices.
Acknowledgements.
This work was partially supported by Independent Research Fund Denmark under Grant 9041-00368B. All of the computations for this project were performed on the UCloud interactive HPC system. Dedicated to the memory of José Apolinário Nunes $\dagger$.